Solve for the equation 5/2t= -3/5 I get 2=t is this true?
- 1/7z - 5 =0 I get z=35 is this right?

Answer by wuwei96815(245)

5/2t = -3/5
t = -3/5 / 5/2 = -3/5 x 2/5 = -6/25
I get t = -6/25
Check this answer by substituting it for t in the original equation.

-1/7z -5 = 0
-1/7z = 5
z = 5 / -1/7 = 5 x -7/1 = -35
I get z = -35
Check this answer by substituting it for z in the original equation.